Please help me solve this word problem. The permimeter of a rectangle is 76cm. The length is 17cm more than the width. Find the length and the width.
Answer by ankor@dixie-net.com(22740)

Please help me solve this word problem. The perimeter of a rectangle is 76cm.
The length is 17cm more than the width. Find the length and the width.
:
The perimeter:
2L + 2W = 76
:
It says,"The length is 17cm more than the width"; therefore:
L = W+17
:
Replace L with (W+17) in the 1st equation, find W, when you can find L
2(W+17) + W = 76
:
Check your values for L & W by finding the perimeter.